A deposit of $4,000 earns interest at 12% compounded semiannually for the first two years, and then 9% compounded monthly thereafter. How much will be in the account five years after the deposit was made? ____________

 


1
Expert's answer
2021-02-16T09:12:31-0500

Solution.

The formula for calculating compound interest:


"B=A(1+\frac{p}{100})^n,"


where "B" is the future value;

"A" - current value;

"p" - interest rate for the settlement period (day, month, year, ...);

"n" is the number of settlement periods.

Сalculate the amount of the deposit in 2 years:

"B=4000(1+\frac{12}{100})^4=\newline
4000*1.12^4=6294.08"$

Сalculate the amount of the deposit in the next 3 years:

"B=6294.08(1+\frac{9}{100})^{36}=\newline
6294.08*1.09^{36}=140050.96"$

Answer. "140050.96"$